OpenStudy (anonymous):

help me find int_{0}^{1} x sqrt{x^2+1}dx =

OpenStudy (turingtest):

u-sub\[u=x^2+1\]

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x \sqrt{x^2+1}dx = \]

OpenStudy (anonymous):

i think u=sprt(x^2+1) better

OpenStudy (anonymous):

(a)\[2\sqrt{2}-1/3\] (b)\[2(2\sqrt{2}-1)/3\]

OpenStudy (turingtest):

if\[u=x^2+1\]then\[du=?\]